OpenStudy (anonymous):

For questions 8-10, use the radical equation: radical x-2 = 5 8. squaring both sides will yield a)x – 2 = 25 b)x – 2 = 5 c) x – 4 = 25 d) x2 – 4 = 25 9. The correct solution to the radical equation is: a) 3 b) 21 c) 27 d)29 10. Squaring both sides of the equation to remove the radicand first: a)Is always the best first step b) Can sometimes cause the problem to look more difficult c)Can always make the problem more difficult d) None of the above

OpenStudy (anonymous):

\[\sqrt{x-2}=5\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[(\sqrt{x-2})^2=5^2\]

OpenStudy (anonymous):

x-2=25 x= 25+2 x= 27

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so its a?

OpenStudy (anonymous):

8)a 9)c

OpenStudy (anonymous):

10)a Squaring both sides of the equation to remove the radicand first:

OpenStudy (anonymous):

thank you